Problem 10.13 Question 13 of 20 Check My Work Click here to read the eBook: Cost of New Common Stock, re COST OF COMMON EQUITY WITH FLOTATION Banyan Co.s common stock currently sells for $45.25 per share. The growth rate is a constan·8 and the company has an expected di dena y eld or 5% h expected long-run dividend payout ratio is 35%, and the expected return on equity ROE)is 12% New stock can be sold to the public atte a rent price but a tation cost of 15% would be incurred, what would be the cost of new equity? Round your answer to two decimal places. Do not round your intermediate calculations Hide Feedback Incorrect Check My Work
Answer #1

Cost of new equity = {Dividend/[Price(1-Flotation cost)]}+growth rate

= {2.35492 / [45.25(1-.15)]}+.078

= {2.35492/38.4625}+.078

=.0612+.078

=0.1392= 13.92%

Therefore, answer is 13.92% cost of new equity.

Above, we arrived dividend by the following calculation:

Expected dividend payout= 35% of ROE = 35*12=4.2%

Dividend = (45.25*.05)*(1+.042)=2.26*1.042 = 2.35492